2/X+3-3X+5/X^2+4X+3=5/X+1
I keep getting negetative three but when i plug it in it doesn't work and i dont know how else i can solve it can you PLEASEE help!!!

Answer by stanbon(75887)

2/(X+3)-(3X+5)/(X^2+4X+3) =5/(X+1)
--------
lcm = (x+3)(x+1) which is x^2+4x+3
--------
Multiply thru by the lcm to get:
2(x+1) - (3x+5) = 5(x+3)
2x+2 -3x-5 = 5x+15
6x = -18
x = -3
But x cannot be -3 for that would make the 1st fraction's denominator be zero.
--------
Therefore: there is no solution to the original equation.
-----------
Cheers,
Stan H.